OpenStudy (kobeni-chan):

When solving a radical equation, Beth and Kelly came to two different conclusions. Beth found a solution, while Kelly's solution did not work in the equation. Create and justify two situations: one situation where Beth is correct and a separate situation where Kelly is correct.

OpenStudy (kobeni-chan):

Will medal! I need this done tonight

OpenStudy (perl):

some radical equations have no solution

OpenStudy (perl):

here is an example of a radical equation that has no solution sqrt( x + 3) = -1

OpenStudy (perl):

here is a radical equation that has a solution sqrt( x + 3) = 2

OpenStudy (anonymous):

sqrt(x)=-3 if squared both sides x=9 and if check sqrt(9)=-3 ? 3=-3?!

OpenStudy (perl):

yes thats a simpler example

OpenStudy (perl):

The equation sqrt(x) = 3 does have a solution: if you square both sides you get x = 3^2 =9 Check: sqrt(9) = 3 sqrt(x) = -3 does not have a solution if you square both sides you get x = (-3)^2 = 9 Check: sqrt(9) = 3 , not -3

OpenStudy (perl):

so solving a radical equation by squaring does not necessarily produce a solution. sometimes we get what is called 'extraneous solution'

OpenStudy (kobeni-chan):

ok